# From Restless to Contextual: A Thresholding Bandit Reformulation for Finite-horizon Improvement

### v4

@@ Appendix F Technical Details In Section 4
$\leq\mathbb{E}\left[\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right]+\mathbb{E}_{\pi_{t+1}}\left[V_{t+1}^{\bm{\pi}^{b}}\left(\bm{s}_{t+1}\right)\mid\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]$
$=\mathbb{E}\left[\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right],$
@@
$\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right].$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right]$ $=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random}\right\}\right]$ (F.58)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random}\right\}\right],$ (F.59)
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random}\right\}\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random},\exists(m^{\prime},s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random},s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{\mathrm{random},s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{k=0}^{\infty}\mathbbm{1}\left\{\mathrm{random},N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime},\tau_{k}^{m^{\prime}}(s^{\prime})\leq T\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{\mathrm{random},N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime}\right\}\right]$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random}\right\}\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random},\exists(m^{\prime},s^{\prime})\in\mathcal{G},\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\sum_{(m^{\prime},s^{\prime})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$ (F.61)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m^{\prime}}(s^{\prime})\leq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right].$ (F.62)
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=0}^{\infty}\mathbbm{1}\left\{s_{\tau_{k}^{m}(s)}^{m}=s,\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\},\tau_{k}^{m}(s)\leq T\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$+\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\notin[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$+\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\notin[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}]\right\}\right]+\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{t_{\mathrm{random}}^{k}\leq k_{3}^{k}\right\}\right],$
where $N_{s,m}^{(\tau_{k}^{m}(s))}(r):=\sum_{l=1}^{k}\mathbbm{1}\left\{a_{\tau_{l}^{m}(s)}^{m}=0,s_{\tau_{l}^{m}(s)}^{m}=s,\mathrm{random}\right\}$ denotes the total number of times that action 0 is given to agent $m$ when $m$ is in state $s$ and the algorithm is doing random exploration, $t^{k}_{\mathrm{random}}:=\sum_{l=1}^{m}\sum_{l=1}^{k}\mathbbm{1}\left\{\mathrm{random}\right\}$ denotes the total number of times when agent $m$ is in state $s$ and the algorithm is doing random explorations. We let $k_{1}=\max\left\{\frac{M-B}{2M},\frac{2M-3B}{2M}\right\},k_{2}=\min\left\{\frac{3(M-B)}{2M},\frac{2M-B}{2(M-B)},k_{3}^{k}=\frac{C\log(k)}{2\min\{\Delta^{2},1/2\}}\right\}$
@@
$\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})t_{\mathrm{random}}^{k},\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\geq k_{1}k_{3}^{k},N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})k_{3}^{k},\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbbm{1}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\geq k_{1}k_{3}^{k},N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})k_{3}^{k},\left|\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)-I^{m}(s)\right|\geq\frac{\Delta}{2}\right\}\right]$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\text{not~random}\right\}\right]$
$=\sum_{(m^{\prime},s^{\prime})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbbm{1}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$

### v5

@@ Appendix F Technical Details In Section 4
$\leq\mathbb{E}\left[\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right]+\mathbb{E}_{\pi_{t+1}}\left[V_{t+1}^{\bm{\pi}^{b}}\left(\bm{s}_{t+1}\right)\mid\bm{s}_{t+1},\bm{a}_{t+1},\mathcal{H}_{t}\right]$
$=\mathbb{E}\left[\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=1\right\}\right],$
@@
$\sum_{I^{m}(s)\leq\gamma}\Delta_{s}^{m}\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right].$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s\right\}\right]$ $=\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random}\right\}\right]$ (F.58)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random}\right\}\right],$ (F.59)
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random}\right\}\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random},\exists(m^{\prime},s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{random},s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{\mathrm{random},s_{t}^{m^{\prime}}=s^{\prime},N_{s^{\prime},m^{\prime}}^{(t)}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(t)},\forall m^{\prime\prime}\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{k=0}^{\infty}\mathbb{I}\left\{\mathrm{random},N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime},\tau_{k}^{m^{\prime}}(s^{\prime})\leq T\right\}\right]$
$\leq\sum_{m^{\prime}}\sum_{s^{\prime}}\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{\mathrm{random},N_{s^{\prime},m^{\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))}\leq N_{s_{t}^{m^{\prime\prime}},m^{\prime\prime}}^{(\tau_{k}^{m^{\prime}}(s^{\prime}))},\forall m^{\prime\prime}\right\}\right]$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random}\right\}\right]$
$=\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\mathrm{not~random},\exists(m^{\prime},s^{\prime})\in\mathcal{G},\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\sum_{(m^{\prime},s^{\prime})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$ (F.61)
$+\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m^{\prime}}(s^{\prime})\leq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right].$ (F.62)
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2},s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=0}^{\infty}\mathbb{I}\left\{s_{\tau_{k}^{m}(s)}^{m}=s,\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\},\tau_{k}^{m}(s)\leq T\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$=\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$+\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\notin[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$+\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\notin[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}]\right\}\right]+\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{t_{\mathrm{random}}^{k}\leq k_{3}^{k}\right\}\right],$
where $N_{s,m}^{(\tau_{k}^{m}(s))}(r):=\sum_{l=1}^{k}\mathbb{I}\left\{a_{\tau_{l}^{m}(s)}^{m}=0,s_{\tau_{l}^{m}(s)}^{m}=s,\mathrm{random}\right\}$ denotes the total number of times that action 0 is given to agent $m$ when $m$ is in state $s$ and the algorithm is doing random exploration, $t^{k}_{\mathrm{random}}:=\sum_{l=1}^{m}\sum_{l=1}^{k}\mathbb{I}\left\{\mathrm{random}\right\}$ denotes the total number of times when agent $m$ is in state $s$ and the algorithm is doing random explorations. We let $k_{1}=\max\left\{\frac{M-B}{2M},\frac{2M-3B}{2M}\right\},k_{2}=\min\left\{\frac{3(M-B)}{2M},\frac{2M-B}{2(M-B)},k_{3}^{k}=\frac{C\log(k)}{2\min\{\Delta^{2},1/2\}}\right\}$
@@
$\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\in[k_{1}t^{k}_{\mathrm{random}},k_{2}t^{k}_{\mathrm{random}}],N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})t_{\mathrm{random}}^{k},\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2},t^{k}_{\mathrm{random}}\geq k_{3}^{k}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\geq k_{1}k_{3}^{k},N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})k_{3}^{k},\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)\geq\gamma+\frac{\Delta}{2}\right\}\right]$
$\leq\mathbb{E}\left[\sum_{k=0}^{T}\mathbb{I}\left\{N_{s,m}^{(\tau_{k}^{m}(s))}(r)\geq k_{1}k_{3}^{k},N_{s,m,1}^{(\tau_{k}^{m}(s)}\geq(1-k_{2})k_{3}^{k},\left|\hat{I}_{\tau_{k}^{m}(s)}^{m}(s)-I^{m}(s)\right|\geq\frac{\Delta}{2}\right\}\right]$
@@
$\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\text{not~random}\right\}\right]$
$=\sum_{(m^{\prime},s^{\prime})\in\mathcal{G}}\mathbb{E}\left[\sum_{t=0}^{T}\mathbb{I}\left\{a_{t+1}^{m}=1,s_{t}^{m}=s,\hat{I}_{t}^{m}(s)\geq\hat{I}_{t}^{m^{\prime}}(s^{\prime}),s_{t}^{m^{\prime}}=s^{\prime}\right\}\right]$